OpenStudy (anonymous):

-8=1/3x+x please solve and show the steps thank you

OpenStudy (anonymous):

Multiply each side by 3.\[3(-8)= \left(\frac{4 x}{3}\right)3\]\[-24=4 x \]Divide each side by 4\[\frac{-24}{4}=\frac{4 x}{4}\]-6 = x x = -6